# AI Safety in the Eyes of the Downstream Developer: A First Look at Concerns, Practices, and Challenges

### v1

@@ 4.3. RQ3: Perceived Challenges
Figure 5. Survey results on agreement on AI safety challenges.
@@
Theme 1: Infrastructure gaps refers to the absence of artefacts and tools for managing AI safety concerns. Evaluation and benchmark gaps highlights the lack of metrics and benchmarks for evaluating AI safety. P4 noted: “There is no metric to decide if the model is safe enough.”. The lack of benchmark includes both an absence of standardised datasets for specific AI safety concerns and limited datasets marginalised groups such as less common cultures. Similarly, monitoring system gaps reflect the absence of systems to track deployed AI software. For instance, P13 emphasised the need for a dashboard-like monitoring system to proactively detect safety breaches over relying on user feedback. Our survey confirms these challenges, receiving 85% agreement on the challenges of evaluation and benchmark gaps, as well as monitoring system gaps.
@@
Theme 2: Technical and model-related difficulties includes both inherent and external challenges in addressing AI safety issues within model and software systems. Model interpretability and transparency issue hinders model understanding, making safety-oriented fine-tuning and testing difficult. As P12 noted: “AI models are often black boxes, making explainability difficult, which complicates troubleshooting unexpected behaviour.” This was the most acknowledged challenge e.g., with 87% agreement in the survey. Performance safety tradeoffs arises from accuracy loss in safety fine-tuning and efficiency reductions from behavioural guardrails, complicating the balance between performance and safety. This challenge was recognised by 64% of survey participants, though 14% disagreed. The key challenge in model monitoring and maintenance issues is tracking behavioural drift over time and updating systems for AI safety concerns. As P16 noted: “So the questions that are asked, the issues that are faced these change overtime. And so what’s difficult, especially for a customer facing issue is the ability to continue to update the information”.This challenge is worsened by insufficient monitoring metrics and the lack of concrete thresholds for necessary updates. Additionally, adapting updates is difficult due to the challenge of reproducing the observed safety issues. 83% of our survey participants agreed with this challenge.
@@
Theme 3: Lack of mature process refers to insufficient and unreliable methods for addressing AI safety concerns during development. Safety policy interpretation gap highlights the lack of well-defined policies and systematic approaches. For instance, P9 emphasised ambiguity in policy interpretation: “For example, you cannot use untrusted third-party library. The definition is not really clear … so sometimes we need to decide by ourselves.”. P12 noted the complexity in ensuring compliance with legal and ethical standards in multi-regulatory environments. This gap extends to the absence of standardised approaches for managing various AI safety concerns, with P18 adding: “there’s a lack of clarity or even in my opinion, community-agreed best practices on how to do monitoring”. In our survey, 79% respondents agreed that safety policy interpretation gaps pose a challenge. Unreliable methodology refers to the flaws in methodologies themselves, encompassing the limitations of data annotation and crowdsourcing processes (Hsueh et al., 2009). Further concerns include the questionable reliability of user feedback and the inherent difficulty of comprehensively anticipating AI safety issues during development. This challenge was also strongly supported, with 84% of survey participants acknowledging it as a difficulty.
@@
Theme 4: Resources constraints refers to the lack of essential resources. The main subtheme is lack of financial and time resources, as ensuring AI safety requires extra steps, specialised expertise, and more time, creating a significant financial burden. P14 noted: “since we are a small company, we don’t have a big budget”. This challenge was supported by 74% of the survey respondents. With these additional steps for handle AI safety issues also comes lack of human resources. The process of high quality dataset curation requires human resources. Also, most AI-based software developers are not specialised in handling AI safety issues, necessitating the involvement of more experts, as highlighted by P15: “But you don’t have any specialist for the AI safety”. This challenge was agreed by 83% of participants and received the highest score of 4.26, identified as one of the biggest challenges developers face.
@@
Theme 5: Barriers to technical understanding and information access encompasses challenges in accessing and understanding AI safety implementation. The subtheme technology knowledge gaps highlights difficulties arising from unfamiliarity with AI safety terminology and techniques. P2 noted the challenge of technical literacy when discussing AI model privacy metrics: “For our general users, I think most people they do not know the concept of privacy budgets”. This issue also affects cross-team collaboration due to inconsistencies in terminology among stakeholders (Nahar et al., 2022). This challenge is agreed by 86% of the survey participants. Meanwhile, poor documentation refers to incomplete or absent safety evaluation records for PTMs and datasets. P18 highlighted a critical mismatch between available and required safety information: “There’s often a misalignment between what benchmarks are publicly available for the particular task and domains that are actually relevant to the project.” This challenge received the highest average score (4.26), indicating a significant lack of documentation efforts.
@@ 5. Implications
AI-based software developers. Implications are discussed based on the development process (Amershi et al., 2019). (1) Preparation and PTM Selection stage: Based on practices in Table 3, interview participants rarely reflected on deeper understanding and assessment for safety concerns. These actions mainly occur in initial stages of development processes, suggesting insufficient attention from developers in analysing and anticipating AI safety risks from the outset. Meanwhile, PTM attributes and potential risks have been highlighted as essential considerations at the initial stage (Jiang et al., 2023). Therefore, AI-based software developers should dedicate more efforts to assessing risks and intended usage early on, as this proactive approach can reduce downstream challenges and help prevent common misuse and exploitation failures observed in the AI incident database.
@@
(2) Model Development and Testing: Developers frequently adopted practices related to model development and testing, especially for “test and evaluate AI model safety” and “improve safety assurance by iteratively refining models and systems”, which are found even more frequently when dealing with model behaviour concerns in AI safety. However, frequent failures in the AI incident database raise questions about their effectiveness. In fact, we observed varying levels of systematic and comprehensive approaches, with some participants testing AI safety by randomly probing a few examples. Others invested more effort, evaluating real-time data using a combination of automated metrics and human assessment. Recent machine learning research has introduced methods to assess bias (Salewski et al., 2023), toxicity contents (Shaikh et al., 2022), reliability (Wang et al., 2023), with various treatments (Jung et al., 2024; Chen and Liu, 2023). However, knowledge gaps, insufficient documentation, and lack of policies hinder developers from effectively leveraging these methods. To mitigate this knowledge transfer gap, downstream developers should actively seek out new techniques. Given the rapid pace of advancements, aggregator platforms like Reddit or Twitter (Aniche et al., 2018) can serve as valuable resources for staying updated on emerging AI safety solutions. However, developers should also be mindful as not all sources on the Internet are trustworthy, which could cause backfire if adopt blindly.
@@
(3) Model Deployment and Maintenance: Similar to the previous phase, downstream developers mentioned varying practices, though fewer compared to the previous stage were reflected in interviews and the survey (see Table 3 and Figure 4), such as monitor input and output and collect user feedback. This phase is particularly challenging due to limited tool support, unpredictable user interactions, and performance drift. Given the frequent failures for misuse and exploitation in the AI incident database, more frequent and careful practices in this stage are needed. The knowledge transfer gap remains for developers, as AI safety monitoring measures are proposed (Lee et al., 2025), catering for unseen scenarios. Downstream developers still need to more proactively engage in acquiring advancements, by reading papers or using aggregator platforms.
@@
Despite efforts to apply AI safety frameworks in regulatory practices (nis, [n. d.]; eur, [n. d.]) and curate question banks for responsible AI risk assessment (Lee et al., 2024), these initiatives emphasise principles over concrete development actions at each stage. An evolving checklist outlining specific actions for addressing various AI safety concerns at different stages of AI software development could help standardise practices and serve as a guideline for developers to mitigate these issues effectively (Braz et al., 2022), particularly in small companies and open-source projects, where safety norms are often lacking. Moreover, drawing an analogy to common criteria evaluation assurance level (EAL, [n. d.]), policy bodies could implement an automated assessment framework to evaluate AI systems across different safety dimensions or based on the aforementioned checklist. AI-based software meeting necessary safety requirements could receive tiered certification indicating its compliance level with specific safety requirements, comparable to ISO 42001 certification (iso, [n. d.]) but lightweight, tiered, and automated. This would help downstream users adopt AI-based software that meet established safety standards. Stronger regulations are also needed to prevent misuse. A frequent issue observed in the AI incident database is that many AI-based software applications are used outside their intended purposes, leading to harmful consequences such as employees being fired solely based on AI decisions or non-medical certified software used for medical purposes.
@@
Researchers. For software engineering researchers, our AI incident database analysis reveals model reliability failures constitute more than half of the reported cases, while failures caused by irresponsible, unintended use or over-reliance of the delivered software is also observed taking 30.9% of the incidents. Furthermore, 15.1% of incidents involve both types of failures, highlighting the widespread premature adoption of unreliable AI-based software under unintended scenarios. Meanwhile, numerous research efforts have proposed solutions for AI safety concerns (Salewski et al., 2023; Chen and Liu, 2023), alongside evolving regulations on AI safety (iso, [n. d.]; eur, [n. d.]). However, with the rapid emergence of new research techniques and policy adjustments, downstream developers often struggle to keep pace, leading to information overload. The frequent occurrence of model behaviour failures and irresponsible usage, coupled with the abundant research and regulations, underscores a significant knowledge transfer gap between AI safety research and practical implementation.
@@
Therefore, research aimed at understanding the challenges that contribute to this knowledge transfer gap between upstream AI researchers and regulators and downstream developers can help bridge this divide. Similar to studies on DevOps education challenges (Fernandes et al., 2022), identifying barriers in AI safety knowledge dissemination could improve its practical adoption in development. Regarding automated tool supports, developing recommender systems or summarisation techniques that process up-to-date policy and research documentation could be a promising direction. For example, TaskNavigator, proposed by Treude et al. (Treude et al., 2014), helps developers navigate software documentation based on different programming tasks. Similarly, such systems could provide targeted policy guidelines and highlight the most relevant evaluation and mitigation solutions based on users’ queries about specific AI concerns at different stages of development. This approach could help standardise AI safety practices in AI-based software development.
@@ 6. Threats to Validity
Construct validity: AI safety definitions vary in scope, from direct harm (Amodei et al., 2016) to broader use case concerns (Hernández-Orallo et al., 2020). To capture a comprehensive perspective, we asked developers about their concerns rather than imposing a predefined definition, minimising response bias. For development processes, we referred to several AI-based software development processes (Amershi et al., 2019; Jiang et al., 2023) and synthesised to the context of PTM. The development process is then used to structure our interview protocol, with refinement conducted based on pilot study feedbacks, facilitating systematic investigation for each RQ.
@@
Internal validity: One potential limitation is the relatively small number of interview participants. However, we reached subtheme-level saturation, and no new subtheme emerged with five additional participants. For the survey participants, we have to rely on their self-claimed identity. We mitigate the risk by following the process by Schmidt et al. (Schmidt et al., 2023) to filter out participants whose answer in their developed AI-based software and job roles mismatch to ensure the quality of our survey responses. Another threat lies in the subjectivity of coding and synthesising process. To mitigate this, we conducted a rigorous examination of the transcripts, with three authors reviewing the coding assignments and aggregations. The final themes and subthemes were established through group consensus. Additionally, since interview participants did not necessarily develop the AI systems in the incident database, we did not establish direct links but provided directional discussions.
@@
External validity: First, we recruited participants from Hugging Face in descending order of organisation size, which may over-represent large companies. However, the participants’ current roles involve diverse organisation sizes as indicated by the demographics. Second, we did not filter interview candidates based on the type of models or tasks they develop. As indicated by participants demographics, our results lean more towards NLP-based software developers. Last, given the complexity of AI-based software development, each participant may not reflect in all development stages. To mitigate this, we continued recruitment until saturation was reached across all subthemes.

### v3

@@ 4.3. RQ3: Perceived Challenges
Figure 6. Survey results on agreement on AI safety challenges. The full details about the distribution is in our replication package section 8
@@
Theme 1: Infrastructure gaps refers to the absence of artefacts and tools for managing AI safety concerns. Evaluation and benchmark gaps highlights the lack of metrics and benchmarks for evaluating AI safety. P4 noted: “There is no metric to decide if the model is safe enough”. The lack of benchmark includes both an absence of standardised datasets for specific AI safety concerns and limited datasets marginalised groups such as less common cultures. Similarly, monitoring system gaps reflect the absence of systems to track deployed AI software. For instance, P13 emphasised the need for a dashboard-like monitoring system to proactively detect safety breaches over relying on user feedback. Our survey confirms these challenges, receiving 85% agreement on the challenges of evaluation and benchmark gaps, as well as monitoring system gaps.
@@
Theme 2: Technical and model-related difficulties includes both inherent and external challenges in addressing AI safety issues within model and software systems. Model interpretability and transparency issue hinders model understanding, making safety-oriented fine-tuning and testing difficult. As P12 noted: “AI models are often black boxes, making explainability difficult, which complicates troubleshooting unexpected behaviour”. This was the most acknowledged challenge e.g., with 87% agreement in the survey. Performance safety tradeoffs arises from accuracy loss in safety fine-tuning and efficiency reductions from behavioural guardrails, complicating the balance between performance and safety. This challenge was recognised by 64% of survey participants, though 14% disagreed. The key challenge in model monitoring and maintenance issues is tracking behavioural drift over time and updating systems for AI safety concerns. As P16 noted: “So the questions that are asked, the issues that are faced these change overtime. And so what’s difficult, especially for a customer facing issue is the ability to continue to update the information”. This challenge is worsened by insufficient monitoring metrics and the lack of concrete thresholds for necessary updates. Additionally, adapting updates is difficult due to the challenge of reproducing the observed safety issues. 83% of our survey participants agreed with this challenge.
@@
Theme 3: Lack of mature process refers to insufficient and unreliable methods for addressing AI safety concerns during development. Safety policy interpretation gap highlights the lack of well-defined policies and systematic approaches. For instance, P9 emphasised ambiguity in policy interpretation: “For example, you cannot use untrusted third-party library. The definition is not really clear … so sometimes we need to decide by ourselves”. P12 noted the complexity in ensuring compliance with legal and ethical standards in multi-regulatory environments. This gap extends to the absence of standardised approaches for managing various AI safety concerns, with P18 adding: “there’s a lack of clarity or even in my opinion, community-agreed best practices on how to do monitoring”. In our survey, 79% respondents agreed that safety policy interpretation gaps pose a challenge. Unreliable methodology refers to the flaws in methodologies themselves, encompassing the limitations of data annotation and crowdsourcing processes (Hsueh et al., 2009). Further concerns include the questionable reliability of user feedback and the inherent difficulty of comprehensively anticipating AI safety issues during development. This challenge was also strongly supported, with 84% of survey participants acknowledging it as a difficulty.
@@
Theme 4: Resources constraints refers to the lack of essential resources. The main subtheme is lack of financial and time resources, as ensuring AI safety requires extra steps, specialised expertise, and more time, creating a significant financial burden. P14 noted: “since we are a small company, we don’t have a big budget”. This challenge was supported by 74% of the survey respondents. With these additional steps for handle AI safety issues also comes lack of human resources. The process of high quality dataset curation requires human resources. Also, most developers of AI-based software are not specialised in handling AI safety issues, necessitating the involvement of more experts, as highlighted by P15: “But you don’t have any specialist for the AI safety”. This challenge was agreed by 83% of participants and received the highest score of 4.26, identified as one of the biggest challenges developers face.
@@
Theme 5: Barriers to technical understanding and information access encompasses challenges in accessing and understanding AI safety implementation. The subtheme technology knowledge gaps highlights difficulties arising from unfamiliarity with AI safety terminology and techniques. P2 noted the challenge of technical literacy when discussing AI model privacy metrics: “For our general users, I think most people they do not know the concept of privacy budgets”. This issue also affects cross-team collaboration due to inconsistencies in terminology among stakeholders (Nahar et al., 2022). This challenge is agreed by 86% of the survey participants. Meanwhile, poor documentation refers to incomplete or absent safety evaluation records for PTMs and datasets. P18 highlighted a critical mismatch between available and required safety information: “There’s often a misalignment between what benchmarks are publicly available for the particular task and domains that are actually relevant to the project”. This challenge received the highest average score (4.26), indicating a significant lack of documentation efforts.
@@ 5. Implications
Figure 7. Practices and challenges across development stages.
Developers of AI-based software. Implications are discussed following the ML development process described by Amershi et al. (Amershi et al., 2019) and Jiang et al.  (Jiang et al., 2023). During the coding process for the interview data, we kept notes on the development stage in which each practice and challenge was discussed. Figure 7 illustrates these stages and shows how each subtheme of practices and challenges was allocated. As indicated in section 3.1, we conducted the interviews based on the development process of model selection, model development and testing, and model deployment and maintenance. The model development and testing were further divided to finer granularity of tasks, including data preparation, model training and adjustment, and model evaluation, as shown in the figure.
@@
(1) Preparation and PTM Selection stage: Based on practices in Table 4, interview participants rarely reflected on deeper understanding and assessment for safety concerns. These actions mainly occur in initial stages of development processes as shown in Figure 7, suggesting insufficient attention from developers in analysing and anticipating AI safety risks from the outset. Meanwhile, PTM attributes and potential risks have been highlighted as essential considerations at the initial stage (Jiang et al., 2023). Therefore, developers of AI-based software should dedicate more efforts to assessing risks and intended usage early on, as this proactive approach can reduce downstream challenges and help prevent common misuse and exploitation failures observed in the AI incident database. One possible way to foster the early-stage reflection is through fictional design approaches. For example, fictional research abstract is found to help in anticipatory reflection on AI ethics (Jääskeläinen et al., 2025), and such methods could potentially be adapted to encourage safety-focused reflection during the PTM selection stage.
@@
(2) Model Development and Testing: Developers frequently adopted practices related to model development and testing, especially for the subthemes of test and evaluate AI model safety and improve safety assurance by iteratively refining models and systems as indicated in Figure 7, which are found even more frequently when dealing with model behaviour concerns in AI safety (see Figure 5). However, frequent failures in the AI incident database raise questions about their effectiveness. In fact, we observed varying levels of systematic and comprehensive approaches, with some participants testing AI safety by randomly probing a few examples. Others invested more effort, evaluating real-time data using a combination of automated metrics and human assessment. Recent machine learning research has introduced methods to assess bias (Salewski et al., 2023), toxicity contents (Shaikh et al., 2022), reliability (Wang et al., 2023), with various treatments (Jung et al., 2024; Chen and Liu, 2023). However, knowledge gaps, insufficient documentation, and lack of policies hinder developers from effectively leveraging these methods. To mitigate this knowledge transfer gap, downstream developers should actively seek out new techniques. Given the rapid pace of advancements, aggregator platforms like Reddit or Twitter (Aniche et al., 2018) can serve as valuable resources for staying updated on emerging AI safety solutions. However, developers should also be mindful as not all sources on the Internet are trustworthy, which could cause backfire if adopt blindly.
@@
(3) Model Deployment and Maintenance: Similar to the previous phase, downstream developers mentioned varying practices, though fewer compared to the previous stage were reflected in interviews and the survey (see Table 4 and Figure 5), such as monitor input and output and collect user feedback. This phase is particularly challenging due to limited tool support, unpredictable user interactions, and performance drift. Given the frequent failures for misuse and exploitation in the AI incident database, more frequent and careful practices in this stage are needed. The knowledge transfer gap remains for developers, as AI safety monitoring measures are proposed (Lee et al., 2025), catering for unseen scenarios. Although it may seem unlikely that practitioners routinely consult academic literature, recent work has shown that many PTM reusers actively engage with ML papers to guide their development decisions (Jiang et al., 2023). Thus, encouraging access to curated research insights—whether through papers or aggregator platforms—remains a promising direction for supporting safer deployment practices.
@@
Despite efforts to apply AI safety frameworks in regulatory practices (nis, [n. d.]; eur, [n. d.]) and curate question banks for responsible AI risk assessment (Lee et al., 2024), these initiatives emphasise principles over concrete development actions at each stage. While principles play a crucial role when dealing with abstract and complex parts of the engineering work, their effective implementation often requires translation into concrete actions. In many engineering fields, including software engineering, this can translate to structured artefacts such as checklist (Garousi and Mäntylä, 2016; Ngadiman et al., 2022). An evolving checklist outlining specific actions for addressing various AI safety concerns at different stages of AI software development could help standardise practices and serve as a guideline for developers to mitigate these issues effectively (Braz et al., 2022), particularly in small companies and open-source projects, where safety norms are often lacking. Some AI ethics checklists, such as the Deon Ethics Checklist (dri, [n. d.]), have been investigated in previous studies (Wong et al., 2023), but were found not to be easily accessible or practically usable by downstream developers, and there is still no concrete checklist for AI safety. Incorporating the idea of an AI safety checklist, along with the aforementioned information aggregator on useful strategies, would be helpful for downstream developers to navigate AI safety issues during PTM-reuse software development. Moreover, drawing an analogy to common criteria evaluation assurance level (EAL, [n. d.]), policy bodies could implement an automated assessment framework to evaluate AI systems across different safety dimensions or based on the aforementioned checklist. AI-based software meeting necessary safety requirements could receive tiered certification indicating its compliance level with specific safety requirements, comparable to ISO 42001 certification (iso, [n. d.]) but lightweight, tiered, and automated. This would help downstream users adopt AI-based software that meet established safety standards. Stronger regulations are also needed to prevent misuse. A frequent issue observed in the AI incident database is that many AI-based software applications are used outside their intended purposes, leading to harmful consequences such as employees being fired solely based on AI decisions or non-medical certified software used for medical purposes.
@@
Researchers. For software engineering researchers, our findings point to a significant knowledge transfer gap between AI safety research, policy and real-world implementation. In particular, 86% of the survey participants agree that knowledge gap constitute a challenge for them to incorporate AI safety during the development (see Figure 6). Moreover, safety policy processing gap is consistently identified as a challenge across all the development stages (see Figure 7). Our analysis of the AI incident database reveals model reliability failures constitute more than half of the reported cases, while failures caused by irresponsible, unintended use or over-reliance of the delivered software is also observed taking 30.9% of the incidents. Furthermore, 15.1% of incidents involve both types of failures, highlighting the widespread premature adoption of unreliable AI-based software under unintended scenarios (see Table 4). Meanwhile, numerous research efforts have proposed solutions for AI safety concerns (Salewski et al., 2023; Chen and Liu, 2023), alongside evolving regulations on AI safety (iso, [n. d.]; eur, [n. d.]). Some failures could have been prevented if relevant research were applied. For example, the over-reliance on AI without human oversight, such as cases involving employee layoffs solely based on AI decisions, could have been avoided or at least mitigated by adhering to human-in-the-loop requirements outlined in ISO 42001 (iso, [n. d.]). However, with the rapid emergence of new research techniques and policy adjustments, downstream developers often struggle to keep pace, leading to information overload. The frequent occurrence of model behaviour failures and irresponsible usage, coupled with the abundant research and regulations, underscores a significant knowledge transfer gap between AI safety research and practical implementation.
@@
Therefore, research aimed at understanding the knowledge transfer gap between upstream AI researchers and regulators and downstream developers, and dissemination channels that can help bridge this divide would be valuable. Challenges related to knowledge dissemination should be investigated, as seen in prior studies on software engineering education that highlight difficulties in delivering relevant information to developers (Fernandes et al., 2022). Similarly, just as prior work has explored how failure stories can guide novice developers (Anandayuvaraj et al., 2023), investigating what types of AI safety or failure-related information are most helpful for downstream developers could inform the design of practical tools and interventions. Regarding automated tool supports, developing recommender systems or summarisation techniques that process up-to-date policy and research documentation could be a promising direction. For example, TaskNavigator, proposed by Treude et al. (Treude et al., 2014), helps developers navigate software documentation based on different programming tasks. Similarly, such systems could provide targeted policy guidelines and highlight the most relevant evaluation and mitigation solutions based on users’ queries about specific AI concerns at different stages of development. This approach could help standardise AI safety practices in the development of AI-based software.
For researchers, the practices and challenges we identified also have implications in the context of AI agentic systems. AI agentic systems that decompose complex tasks into smaller subgoals (Khot et al., 2022) and methodically explore different solution paths (Yao et al., 2023) have attracted increasing attention as large language models advance. Although our study did not directly examine the safe development of agentic systems, we suggest that many of the challenges we identified are relevant to the development of such systems. For example, each subtask relying on a single model may face issues such as poor documentation, difficulty in monitoring and maintenance, and a lack of detailed, well-suited benchmark evaluations. However, these challenges could be amplified when multiple components are combined. For instance, iterative improvements to one model could inadvertently impact the behaviour of another, raising additional concerns about the interoperability and coordination of subtasks. This boundary erosion and entanglement of ML models are identified as hidden technical debt in machine learning systems (Sculley et al., 2015). Additionally, Model Context Protocol (MCP) (Hou et al., 2025) emerged as a way for LLMs to interact with external tool. The engineering practices in the interface design and carried information will impact the overall safety of the agentic AI systems. Therefore, unique practices and challenges specific to agentic systems should be investigated in future studies.
@@ 6. Threats to Validity
Construct validity: Construct validity refers to how well a test or assessment measures the specific concept or characteristic it is designed to evaluate. The first construct is the definition of AI safety we followed in the study design. This term lacks a universally agreed scope. Some more restrictive perspectives exclude non-accidental risks and classify them as security over safety. Our study adopts a broader view from Hernandez-Oralloet al. (Hernández-Orallo et al., 2020) that includes both accidental and non-accidental risks. As described in section 3, we did not impose a predefined-definition of AI safety on participants, but only steered the interview into this direction when needed. Therefore, the interpretation of our results should be considered in the context of both accidental and non-accidental risks. Meanwhile, broader implications of indirect harms, such as carbon emissions, are explicitly excluded from the study.
The second construct is our division of the AI-based software development process involving PTM reuse into three stages: (1) preparation and PTM selection, (2) model development and testing, and (3) model deployment and maintenance. This structure is derived from several development processes of AI-based software (Amershi et al., 2019; Jiang et al., 2023), and refined based on pilot study feedbacks to balance response quality and interview duration. While this three-stage framework provides a practical structure, finer granularity is possible. For example, the model development and testing stage could be further divided into data preparation, model training and adjustment, and model evaluation. Conducting interviews at this more detailed level could potentially yield more fine-grained insights.
@@
Internal validity: Internal validity refers to the extent to which the observed results represent the truth in the population studied, and not due to methodological errors. For the survey participants, we have to rely on their self-claimed identity. We mitigate the risk by following the process by Schmidt et al. (Schmidt et al., 2023) to filter out participants whose answer in their developed AI-based software and job roles mismatch to ensure the quality of our survey responses. Another threat lies in the subjectivity of coding and synthesising process. To mitigate this, we conducted a rigorous examination of the transcripts, with three authors reviewing the coding assignments and aggregations. The final themes and subthemes were established through group consensus. Additionally, since interview participants did not necessarily develop the AI systems in the incident database, we did not establish direct links but provided directional discussions.
Additional threats to validity lie in the three data sources. We recruited five of the 16 interview participants from our professional networks, which could introduce selection bias. The downstream developers’ practices are self-reported rather than directly observed in real-world settings. However, the interview sessions were structured by development stage, with questions about AI safety concerns asked first and questions about practices asked afterwards, to help ensure that self-reported practices reflected actual experiences as closely as possible. There are also differences in demographics and system criticality across the data sources. In general, interview participants had more experience with deep learning and similar years of experience in software engineering compared to survey participants, while the AI incident database contained a higher proportion of safety-critical systems. Finally, the cases in the AI incident database were collected and reported by domain experts, which may not be comprehensive and could be subject to selection bias (Anandayuvaraj et al., 2024).
@@
External validity: External validity refers to the extent to which the findings of a study can be generalised to other settings. The first potential limitation is the relatively small number of interview participants. However, we reached subtheme-level saturation, and no new subtheme emerged with five additional participants. Second, we recruited participants from Hugging Face in descending order of organisation size, which may over-represent large companies. However, the participants’ current roles involve diverse organisation sizes as indicated by the demographics. Third, we did not filter interview candidates based on the type of models or tasks they develop. As indicated by participants’ demographics in Table 2, our results lean more towards NLP-based software developers. Last, given the complexity of the development of AI-based software, each participant may not have experience in in all development stages. To mitigate this, we continued recruitment until saturation was reached across all subthemes